A 2.50 g sugar cube (sucrose: C12H22O11) is dissolved in a 350 mL teapot containing 80∘ C water (density of water at 80∘ C= 0.975 g/mL). What is the molality of the sugar solution?
Express the molality with the appropriate units.

I entered these and they are not correct, can you help?
ANSWER 1: Deduction: -3%
0.03molkg
ANSWER 2: Deduction: -3%
343.77gmol
Enter your answer using dimensions of concentration.
1
Expert's answer
2015-11-05T12:43:54-0500
b = n (of solute) / m (of solvent)
solvent – H2O
solute – C12H22O11
ρ (density) = m / V (mass / volume),
so m = ρ * V
m (H2O) = 0.975 (g/mL) * 350 mL = 341.25 g
n = m / M
n (C12H22O11) = 2.5 g / 342 g/mol = 0.0073 mol
b (solution) = 0.0073 mol / 0.3413 g = 0.021 mol/kg
Answer: 0.021 mol/kg